# Can Deception Detection Go Deeper? Dataset, Evaluation, and Benchmark for Deception Reasoning

### v2

@@ Abstract
Deception detection has attracted increasing attention due to its importance in real-world scenarios. Its main goal is to detect deceptive behaviors from multimodal clues such as gestures, facial expressions, prosody, etc. However, these bases are usually subjective and related to personal habits. Therefore, we extend deception detection to deception reasoning, further providing objective evidence to support subjective judgment. Specifically, we provide potential lies and basic facts and then analyze why this sentence may be a lie by combining factual inconsistencies and intent behind them. Compared with deception detection, this task is more applicable to real-world scenarios. For example, in interrogation, the police should judge whether a person is lying based on solid evidence. This paper presents our initial attempts at this task, including constructing a dataset and defining evaluation metrics. Meanwhile, this task can serve as a benchmark for evaluating the complex reasoning capability of large language models. Code and data will be made publicly available.
@@
Kang Chen1,3∗, Zheng Lian1thanks: Equal Contribution,thanks: Corresponding Author, Haiyang Sun1,2, Bin Liu1,2, Jianhua Tao4,5 1Institute of Automation, Chinese Academy of Sciences 2School of Artificial Intelligence, University of Chinese Academy of Sciences 3Peking University, 4Department of Automation, Tsinghua University 5Beijing National Research Center for Information Science and Technology, Tsinghua University {2201210163@stu.pku.edu.cn, lianzheng2016@ia.ac.cn}
@@ 1 Introduction
In this task, our main goal is not to improve the authenticity of deception but to focus on the rationality of reasoning. Therefore, to reduce the cost of data collection, we use GPT-4 to synthesize dialogues with deceptive behaviors. Besides datasets, we define four metrics to comprehensively evaluate the reasoning results: accuracy, completeness, logic, and depth. The main contributions of this paper are summarized as follows:
@@
  • To facilitate subsequent research, we construct a dataset and evaluation metrics for this task.
@@
The rest is organized as follows: Section 2 reviews some recent works. In Section 3, we introduce our data generation pipeline. In Section 4, we define evaluation metrics and report the performance of various LLMs on deception reasoning. Finally, we conclude this paper in Section 5.
@@ 3 Data Generation
In deception reasoning, we pick a potential lie and analyze why this sentence may be a lie by considering factual inconsistencies and the intent behind it. In this task, we focus on the rationality of reasoning rather than the authenticity of deceptive behaviors. Therefore, to reduce the cost of dataset collection, we use GPT-4 to synthesize dialogues containing deceptive behaviors. Specifically, we choose one of the most widely used scenarios in previous works, mock crime Derrick et al. (2010); Pérez-Rosas et al. (2014). We ask GPT-4 to simulate the role-playing between a suspect and a police officer. During interrogation, the suspect should deceive the police officer and escape the crime and the police officer should find out the truth and seize evidence.
We first clarify the definition of three important notations: legal instrument, target content, and action. Then, we introduce the data generation process (see Figure 1 for more details). This section mainly uses GPT-3.5 (“gpt-3.5-turbo-0613”) and GPT-4 (“gpt-4-1106-preview”).
@@ 3.5 Mock Interrogation
Model
Cost
(×$\times 10^{-3}\$$)
Accuracy Completeness Logic Depth Sum
ChatGLM2-6B 1.3 4.00 3.56 4.33 3.44 15.33
WizardLM-13B 3.6 5.20 4.87 6.00 4.38 20.45
Baichuan2-13B 2.1 5.24 5.00 6.25 4.62 21.11
ERINE3.5 0.1 5.40 5.00 6.10 5.10 21.60
Llama2-70B 12.4 5.20 5.65 6.65 5.65 23.15
Qwen-14B 2.2 6.00 5.60 6.70 5.20 23.50
Claude3-Haiku 0.9 6.33 5.89 6.89 5.33 24.44
GPT-3.5 4.2 6.00 5.87 6.87 5.75 24.49
ERINE4.0 3.6 6.60 6.30 7.30 5.80 26.00
GLM-4-9B 2.8 6.67 6.44 7.33 6.33 26.77
Gemini-1.5-Pro 0.7 6.11 6.89 7.67 6.56 27.23
Qwen2-7B 1.8 6.56 6.72 7.72 6.39 27.39
Table 3: Deception reasoning performance of different LLMs. We also provide the inference cost per sample.
@@
3.6 Post-filtering and Statistics
In deception reasoning, we pick a potential lie and analyze why this sentence may be a lie by considering factual inconsistencies and the intent behind it. Figure 4 gives an example to illustrate this process. Specifically, we manually select a potential lie that is more representative of humans (in the red box) and generate analysis results. To ensure the corpus quality, we further conduct post-filtering to remove some dialogues that contain unnatural parts.
Totally, we generate 191 dialogues, and their statistics are summarized in Table 2. In this table, we observe that the average number of turns per dialogue is 34.93, which is sufficient for a short interrogation. In Figure 5, we also provide the distribution of target content length, number of actions, and dialogue turns.
Meanwhile, we analyze the cost of data collection. On average, we spend less than \$2 per dialogue. Compared with existing datasets, subject recruitment and data annotation often require a lot of money, and the cost varies from country to country. But in our country, it costs more than \$2 per dialogue. Therefore, this paper provides a cheaper way to collect data.
@@ 4.1 Evaluation Metrics
In deception reasoning, we need to figure out why a sentence may be a lie by considering factual inconsistencies and the intent behind it. To provide a more comprehensive evaluation, we propose four metrics for deception reasoning:
@@ 4.2 Evaluator
Considering that previous works Zheng et al. (2023); Lian et al. (2023) have demonstrated the consistency between GPT-4 and human assessments, this paper directly uses GPT-4 as the evaluator. To test its stability, we run GPT-4 multiple times. We observe relatively small differences between distinct runs, showcasing the stability of GPT-4 in evaluating the deception reasoning ability.
1 point - Very unnatural: The conversation appears very stiff and unnatural, possibly containing numerous grammar errors, incoherent sentences, or content that is completely unrelated to the context. This type of conversation is difficult to understand and gives off a mechanical or robotic feel, lacking the natural fluency of human communication.
2 points - Somewhat unnatural: Although the conversation conveys basic information, it still seems somewhat unnatural. There may be some linguistic or logical inconsistencies that make the conversation lack the smoothness of natural communication. The conversation may occasionally contain content that is unrelated to the context, requiring further improvement to enhance its naturalness.
3 points - Moderately natural: The conversation is somewhat fluent but still has some issues. There may be some lack of coherence in some places, or occasional unnatural expressions. The conversation can generally stay on topic but still has room for improvement to better simulate natural language communication.
4 points - Fairly natural: The conversation is generally fluent and can convey meaning and emotions well. Although there may be occasional minor unnatural aspects, overall, it closely resembles real human dialogue. The conversation is coherent, able to closely follow the topic, and demonstrates good adaptability and understanding.
5 points - Very natural: The conversation is extremely fluent and natural as if it were a real interaction with a person. There are no language or logical inconsistencies throughout the conversation, maintaining consistency and relevance to the topic. The expression is precise, and adaptable, closely simulating human communication habits and emotional responses, giving a very authentic and comfortable feeling.
Table 4: Prompt for evaluating the dialogue naturalness.
@@ 4.3 Main Results
In this section, we evaluate the deception reasoning performance of different LLMs. Specifically, we select mainstream LLMs, such as Llama2-70B Touvron et al. (2023) and WizardLM-13B Xu et al. (2023). Since our dataset is in Chinese, we also select some LLMs that perform well in Chinese, including Qwen-14B Bai et al. (2023), ChatGLM2-6B Du et al. (2021), and Baichuan2-13B Yang et al. (2023). Experimental results are shown in Table 3. We observe that existing LLMs can deal with deception reasoning to some extent, among which Qwen2 performs the best. Meanwhile, we can also see the development of Chinese LLMs. For example, Qwen2 is better than Qwen and ERINE4.0 is better than ERINE3.5. These results demonstrate the progress of LLMs in reasoning ability.
Meanwhile, Table 3 reports the inference cost per sample for different LLMs. For closed-source models provided by OpenAI, Google, etc., we calculate the inference cost based on the number of tokens and the price per token. For open-source models such as GLM-4-9B and Qwen2-7B, we calculate the inference cost based on the model inference time and the daily price of the machine usage. Specifically, we use Azure Standard_NC12s_v3 (equipped with 2 V100 GPUs) based on the pay-as-you-go pricing in December 2023. Although these costs are not accurate due to price changes, they provide a rough estimate of the inference cost. We find that for open-source LLMs, Llama2-70B is expensive due to its large model size and long inference time. For close-source LLMs, Gemini-1.5-Pro is cheaper than GPT-3.5.
@@
4.4 Dialogue Naturalness Evaluation
In this section, we test the naturalness of our synthetic dialogues. Considering that we use GPT-4 to generate dialogues, we choose another powerful LLM Claude3-Haiku for naturalness evaluation. Specifically, we randomly select 10 real dialogues from a dialogue dataset IEMOCAP Busso et al. (2008) and 10 synthetic dialogues from our dataset. We use the prompts in Table 4 to score the naturalness. Experimental results demonstrate that the average score of real dialogue can reach 4.00 and the average score of synthetic dialogue can reach 3.88. These results reflect the naturalness of our synthetic dialogues for deception reasoning.
4.5 Ablation Study
This paper uses a two-stage strategy and GPT-4 for target content and action extraction (see Section 3.3). In this section, we compare the performance between one-stage and two-stage strategies, as well as GPT-3.5 and GPT-4. For target content extraction, we use the target accuracy as the evaluation metric. If the system extracts non-target content from legal instruments, it will have a low score in this metric. For action extraction, we use the action complexity as the metric. If the system cannot accurately realize action decomposition, it will have high action complexity. Therefore, a good model should have high target accuracy and low action complexity. Experimental results of different strategies are shown in Table 5.
@@ 5 Conclusions
This paper extends deception detection to deception reasoning, further providing objective evidence to support subjective judgment. To facilitate subsequent research, we build a dataset, define evaluation metrics, and open-source data and code. In this paper, we reveal the performance of mainstream LLMs in deception reasoning and show the progress of Chinese LLMs in reasoning ability. Meanwhile, we prove the rationality of our dataset construction strategy and the naturalness of our synthetic dialogues. Moreover, this task can also serve as a reasoning benchmark for current LLMs.
@@
There are several limitations that can be addressed in future research. First, our deception dataset relies on GPT-4, which requires API call costs. Therefore, we only select a part of legal instruments from CAIL2018 instead of using the entire dataset. Future research will consider using all legal instruments for dialogue generation. Secondly, this paper evaluates the performance of mainstream LLMs but does not cover all LLMs. In the future, we will expand the evaluation scope. Thirdly, we focus on the rationality of reasoning rather than the authenticity of deceptive behaviors. Therefore, to reduce the cost of data collection, this paper mainly uses synthetic dialogues. In the future, we will also do some experiments on real interrogation dialogues. Fourthly, video generation has become increasingly popular. In the future, we will synthesize multimodal data and expand text-based deception reasoning to multimodal deception reasoning.

### v3

@@ Abstract
Deception detection has attracted increasing attention due to its importance in real-world scenarios. Its main goal is to detect deceptive behaviors from multimodal clues such as gestures, facial expressions, prosody, etc. However, these bases are usually subjective and related to personal habits. Therefore, we extend deception detection to deception reasoning, further providing objective evidence to support subjective judgment. Specifically, we provide potential lies and basic facts and then analyze why this sentence may be a lie by combining factual inconsistencies and intent behind them. Compared with deception detection, this task is more applicable to real-world scenarios. For example, in interrogation, the police should judge whether a person is lying based on solid evidence. This paper presents our initial attempts at this task, including constructing a dataset and defining evaluation metrics. Meanwhile, this task can serve as a benchmark for evaluating the complex reasoning capability of large language models. Our code and data are provided in the supplementary material.
@@
Kang Chen1,2∗, Zheng Lian1thanks: Equal Contribution,thanks: Corresponding Author, Haiyang Sun3, Rui Liu4, Jiangyan Yi1, Bin Liu1, Jianhua Tao5,6 1Institute of Automation, Chinese Academy of Sciences, 2Peking University, 3Shanghai Jiao Tong University, 4Inner Mongolia University 5Department of Automation, Tsinghua University 6Beijing National Research Center for Information Science and Technology, Tsinghua University 2201210163@stu.pku.edu.cn, lianzheng2016@ia.ac.cn
@@ 1 Introduction
In this task, our main goal is not to improve the authenticity of deception but to focus on the reasonableness of reasoning. Therefore, to reduce the cost of data collection, we use GPT-4 to synthesize dialogues with deceptive behaviors. Besides datasets, we define four metrics to comprehensively evaluate the reasoning results: accuracy, completeness, logic, and depth. The main contributions of this paper are summarized as follows:
@@
  • To facilitate subsequent research, we construct a dataset and evaluation metrics.
@@
Section 2 reviews recent works. In Section 3, we introduce our data generation pipeline. In Section 4, we define evaluation metrics and report the performance of various LLMs on deception reasoning. Finally, we conclude this paper in Section 5.
@@ 3 Data Generation
In deception reasoning, we pick a potential lie and analyze why this sentence may be a lie by considering factual inconsistencies and the intent behind it. In this task, we focus on the reasonableness of reasoning rather than the authenticity of deceptive behaviors. Therefore, to reduce the cost of dataset collection, we use GPT-4 to synthesize dialogues containing deceptive behaviors. Specifically, we choose one of the most widely used scenarios in previous works, mock crime Derrick et al. (2010); Pérez-Rosas et al. (2014). We ask GPT-4 to simulate the role-playing between a suspect and a police officer. During interrogation, the suspect should deceive the police officer and escape the crime and the police officer should find out the truth.
We first clarify the definition of three important notations: legal instrument, target content, and action. Then, we introduce the data generation process (see Figure 1 for more details), which mainly relies on GPT-3.5 (“gpt-3.5-turbo-0613”) and GPT-4 (“gpt-4-1106-preview”).
@@ 3.5 Mock Interrogation
Model Automatic Evaluation Results Manual Evaluation Results
Acc. Com. Log. Dep. Sum Acc. Com. Log. Dep. Sum
ChatGLM2-6B 4.00 3.56 4.33 3.44 15.33 5.12 5.00 4.90 4.70 19.72
WizardLM-13B 5.20 4.87 6.00 4.38 20.45 5.61 5.30 5.41 5.14 21.46
Baichuan2-13B 5.24 5.00 6.25 4.62 21.11 4.56 4.36 4.49 4.16 17.57
ERINE3.5 5.40 5.00 6.10 5.10 21.60 5.71 5.81 5.71 5.13 22.36
Qwen-14B 6.00 5.60 6.70 5.20 23.50 5.91 5.80 5.42 5.40 22.53
Claude3-Haiku 6.33 5.89 6.89 5.33 24.44 6.36 6.11 5.94 5.70 24.11
GPT-3.5 6.00 5.87 6.87 5.75 24.49 6.80 6.58 6.53 6.18 26.09
ERINE4.0 6.60 6.30 7.30 5.80 26.00 6.95 6.78 6.99 6.81 27.53
GLM-4-9B 6.67 6.44 7.33 6.33 26.77 7.56 7.54 7.59 7.55 30.24
Gemini-1.5-Pro 6.11 6.89 7.67 6.56 27.23 7.56 7.51 7.37 7.23 29.67
Qwen2-7B 6.56 6.72 7.72 6.39 27.39 7.41 7.49 7.48 7.41 29.79
PCC scores 0.81 0.87 0.80 0.89 0.86 0.81 0.87 0.80 0.89 0.86
Table 3: Main results of different LLMs on four evaluation metrics. We show both automatic and manual evaluation results, and the last row reports the PCC scores between them.
@@
3.6 Complete Pipeline and Data Statistics
[…]
Dataset statistics are summarized in Table 2. We observe that the average number of turns per dialogue is 34.93, which is sufficient for a short interrogation. In Figure 5, we also provide the distribution of target content length, number of actions, and dialogue turns. Meanwhile, we analyze the cost of data collection. On average, we spend less than \$2 per dialogue. Compared with existing datasets, subject recruitment and data annotation often require a lot of money, and the cost varies from country to country. But in our country, it costs more than \$2 per dialogue. Therefore, this paper provides a cheaper way to collect data.
It should be noted that deception reasoning has some similarities with misinformation detection. However, there are also certain differences between them. As shown in Figure 4, misinformation detection is one aspect of deception reasoning. Besides this, deception reasoning also focuses on other aspects, such as ambiguous expressions.
@@ 4.1 Evaluation Metrics
In deception reasoning, we need to figure out why a sentence may be a lie by considering factual inconsistencies and the intent behind it. To provide a more comprehensive evaluation, we propose four metrics for deception reasoning, whose core definitions are provided below:
@@
During the evaluation, we use more detailed definitions for each metric, as well as the meaning of each score for each metric in our prompts. More details can be found in Tables 6$\sim$9.
@@ 4.2 Evaluator
We conduct both automatic and manual evaluations. Considering that researchers Zheng et al. (2023); Lian et al. (2023) have proven the consistency between GPT-4 and human assessments, we directly use GPT-4 for automatic evaluator. Meanwhile, considering that using one LLM to evaluate other LLMs may lead to bias issues and the possibility of overfitting, we further hire 8 annotators and perform manual evaluation. Each annotator is paid \$14 per hour, which is relatively high in our country.
@@ 4.3 Main Results
This section evaluates the deception reasoning performance of different LLMs. Besides mainstream LLMs such as WizardLM-13B Xu et al. (2023), we also select LLMs that perform well in Chinese. During inference, we input basic facts, synthetic dialogue, and potential lies, and ask LLMs to analyze why this sentence might be a lie. In both automatic and manual evaluation, we use the prompts in Tables 6$\sim$9 and experimental results are shown in Table 3. We observe that existing LLMs can deal with deception reasoning to some extent. Meanwhile, we can see the progress of Chinese LLMs in reasoning ability. For example, Qwen2 is better than Qwen and ERINE4.0 is better than ERINE3.5. Furthermore, Table 3 shows the PCC scores between automatic and manual evaluation results. We observe that manual evaluation results have relatively high similarities with automatic evaluation results, proving the reliability of our automatic evaluation strategy.
Table 4 shows the inference cost per sample for each LLM. For closed-source models provided by OpenAI, Google, etc., we calculate the inference cost based on the number of tokens and the price per token. For open-source models such as GLM-4-9B and Qwen2-7B, we calculate the inference cost based on the model inference time and the daily price of the machine usage. Specifically, we use Azure Standard_NC12s_v3 (equipped with 2 V100 GPUs) based on the pay-as-you-go pricing in December 2023. Although these costs are not accurate due to price changes, they provide a rough estimate of the inference cost. We find that for open-source LLMs, large models are often expensive due to their long inference time. For close-source LLMs, Gemini-1.5-Pro is cheaper than GPT-3.5.
Model Cost (×$\times 10^{-3}\$$)
ChatGLM2-6B 1.3
WizardLM-13B 3.6
Baichuan2-13B 2.1
ERINE3.5 0.1
Qwen-14B 2.2
Claude3-Haiku 0.9
GPT-3.5 4.2
ERINE4.0 3.6
GLM-4-9B 2.8
Gemini-1.5-Pro 0.7
Qwen2-7B 1.8
Table 4: Inference cost per sample for different LLMs.
4.4 Dialogue Naturalness Evaluation
To test the naturalness of our synthetic dialogue, we use the prompt in Table 10 and conduct both automatic and manual evaluations.
In the automatic evaluation process, considering that we use GPT-4 to generate dialogues, we choose another Claude3-Haiku for evaluation. Specifically, we randomly select 10 real dialogues from a dialogue dataset IEMOCAP Busso et al. (2008) and 10 synthetic dialogues from our dataset. The average score of real dialogue can reach 4.00 and the average score of synthetic dialogue can reach 3.88.
In the manual evaluation process, we hire eight annotators and ask them to score the naturalness. We observe that the average score of synthetic dialogue can reach 3.70, close to the automatic evaluation results. All these results reflect the naturalness of our synthetic dialogues and the reliability of our automatic evaluation strategy.
4.5 Ablation Study
This paper uses a two-stage strategy and GPT-4 for target content and action extraction (see Section 3.3). In this section, we compare the performance between one-stage and two-stage strategies, as well as GPT-3.5 and GPT-4. During evaluation, we hire one annotator and randomly select 100 samples. For target content extraction, we define a metric called target accuracy. If the system extracts non-target content from legal instruments, it will have a low score in this metric. For action extraction, we define a metric called action complexity. This metric is related to inappropriate action decomposition. Take the complete actions in Table 1 as an example. These actions are well-decomposed. But if we merge two actions into one action, this decomposition process is inappropriate, leading to an increase in action complexity. Therefore, a good model should have high target accuracy and low action complexity. Experimental results of different strategies are shown in Table 5.
@@ 5 Conclusions
This paper extends deception detection to deception reasoning, further providing objective evidence to support subjective judgment. To facilitate subsequent research, we build a dataset, define evaluation metrics, and open-source data and code. Then, we reveal the performance of mainstream LLMs and show the progress of Chinese LLMs in reasoning ability. Meanwhile, we prove the rationality of our dataset construction strategy and the naturalness of our synthetic dialogues. This task can also serve as a reasoning benchmark for current LLMs.
@@
There are several limitations that can be addressed in future research. First, our deception dataset relies on GPT-4, which requires API call costs. Therefore, we only select a part of legal instruments from CAIL2018 instead of using the entire dataset. Future research will consider using all legal instruments for dialogue generation. Secondly, this paper evaluates the performance of mainstream LLMs but does not cover all LLMs. In the future, we will expand the evaluation scope. Thirdly, we focus on the reasonableness of reasoning rather than the authenticity of deceptive behaviors. Therefore, to reduce the cost of data collection, this paper mainly uses synthetic dialogues. In the future, we will also do some experiments on real interrogation dialogues. Fourthly, video generation has become increasingly popular. In the future, we will synthesize multimodal data and expand text-based deception reasoning to multimodal deception reasoning.
@@
Appendix A Metric Calculation
In deception reasoning, we define four evaluation metrics: accuracy, completeness, logic, and depth. In Tables 6$\sim$9, we provide detailed definitions for each metric, as well as the meaning of each score.
Appendix B Dialogue Naturalness Evaluation
We rank the dialogue naturalness using five scores. In Table 10, we provide the meaning of each score.
We provide facts, a dialogue, and a potential lie. Meanwhile, we provide a model’s analysis results of whether this sentence might be a lie. Please evaluate whether the model’s inference aligns with known facts. If an inference is closer to the real situation or facts, it should receive a higher score in this dimension.
0-2 points (Very low accuracy): Most of the arguments do not align with known facts, with only a small part possibly slightly related. Displays a severe misunderstanding of the facts or selective ignorance.
3-4 points (Low accuracy): Some of the arguments align with the facts, but most of the content is still inaccurate or misleading. There is an attempt to use correct facts, but they are handled improperly or misunderstood.
5-6 points (Moderate accuracy): There is some degree of consistency between the arguments and the facts, but there are still noticeable inaccuracies. Displays a basic understanding of the facts, but lacks thorough or detailed consideration.
7-8 points (High accuracy): Most of the arguments align with the facts, with only a few details or aspects showing deviations. Shows a good understanding of the facts and correct application, but there is still room for improvement.
9-10 points (Very accurate): All or almost all of the arguments strictly align with known facts. Accurately and fully understands and applies the facts, with no obvious errors or omissions.
Table 6: Prompt for evaluating the accuracy.
We provide facts, a dialogue, and a potential lie. Meanwhile, we provide a model’s analysis results of whether this sentence might be a lie. Please evaluate whether the model has considered all relevant information and details. A good inference should be comprehensive, without omitting any key points. I will provide a "possible answer" which can be considered a reliable answer scoring 8 or above, as a reference for completeness scoring.
0-2 points (Very low completeness): The inference is extremely one-sided or missing significant content, with almost no resemblance to the "possible answer." Ignores key aspects of the problem, possibly only scratching the surface. Lacks basic understanding of the problem, with results far from the "possible answer."
3-5 points (Moderate completeness): The inference includes some key aspects of the problem but still falls far short of the "possible answer." There is an attempt to address the problem comprehensively, but some important aspects or details are overlooked. The inference has some resemblance to the "possible answer," but there are still clear omissions or misunderstandings.
6-7 points (High completeness): The inference is fairly comprehensive, covering most key aspects and closely aligning with the "possible answer." Able to understand and respond well to the core requirements of the problem, though there may still be some omissions in details. The inference has a high degree of similarity to the "possible answer," but there is still room for improvement.
8 points (Very high completeness): The inference is very comprehensive, covering all key aspects of the problem and highly consistent with the "possible answer." Demonstrates a deep understanding of the problem, responding accurately and thoroughly to all aspects. The result shows depth and detail, with almost no omissions or misunderstandings.
9-10 points (Exceeds completeness): Not only highly consistent with the "possible answer," but also offers innovation or further depth. Provides additional insights.
Table 7: Prompt for evaluating the completeness.
We provide facts, a dialogue, and a potential lie. Meanwhile, we provide a model’s analysis results of whether this sentence might be a lie. Please evaluate whether the inference is logically coherent and well-structured. If the model provides an inference that is disorganized or self-contradictory, it should score lower in this dimension.
0-2 points (Low level): The inference has almost no logical coherence, possibly entirely based on guesswork or conjecture. There is no clear connection between evidence and conclusions, or no evidence is used at all. The reasoning process is chaotic and lacks organization, possibly deviating completely from the core of the problem.
3-5 points (Moderate level): The inference has some logical coherence, but there may be noticeable logical gaps or errors. Some relevant evidence is used, but the application of evidence is either inappropriate or insufficient. The reasoning process, while somewhat structured, may lack in-depth analysis in key areas.
6-8 points (Good level): The inference has strong logical coherence, with few logical gaps. Evidence is used appropriately and sufficiently to support the conclusion. The reasoning process is clear and well-organized, with in-depth analysis of key parts of the problem.
9-10 points (Excellent level): The inference is highly logical, with almost no logical gaps. Evidence is used extremely appropriately and sufficiently, strongly supporting the conclusion. The reasoning process is very clear and well-organized, with in-depth exploration of various aspects of the problem, possibly providing new insights or solutions.
Table 8: Prompt for evaluating the logic.
We provide facts, a dialogue, and a potential lie. Meanwhile, we provide a model’s analysis results of whether this sentence might be a lie. Please evaluate whether there has been deep thought and analysis or if it remains superficial. A high-quality inference should be able to deeply explore underlying reasons and motivations.
0-2 points (Superficial thinking): The inference is superficial, staying only at the surface level. Lacks exploration of underlying reasons and motivations, ignoring the complexity and deeper factors of the problem. The conclusion may be overly simple and direct, not showing multi-angle or in-depth consideration of the problem.
3-5 points (Basic depth): The inference shows some depth, but still not comprehensive or deep enough. While there is an attempt to explore underlying reasons and motivations, the analysis still appears shallow or incomplete. The reasoning process has some logical coherence, but lacks depth and complexity.
6-8 points (Good depth): The inference shows good depth, exploring various aspects of the problem fairly comprehensively. There is in-depth analysis of underlying reasons and motivations, providing deep insights. Although the depth is high, there may still be room for further exploration in some areas.
9-10 points (Extremely deep): The inference is extremely deep and comprehensive, uncovering the core and deeper factors of the problem. The analysis of underlying reasons and motivations is profound, offering unique and deep insights. The reasoning process displays a high level of logical coherence and depth, exploring the problem from multiple angles, showing high-quality thinking.
Table 9: Prompt for evaluating the depth.
1 point: very unnatural. The conversation appears very stiff and unnatural, possibly containing numerous grammar errors, incoherent sentences, or content that is completely unrelated to the context. This type of conversation is difficult to understand and gives off a mechanical or robotic feel, lacking the natural fluency of human communication.
2 points: somewhat unnatural. Although the conversation conveys basic information, it still seems somewhat unnatural. There may be some linguistic or logical inconsistencies that make the conversation lack the smoothness of natural communication. The conversation may occasionally contain content that is unrelated to the context, requiring further improvement to enhance its naturalness.
3 points: moderately natural. The conversation is somewhat fluent but still has some issues. There may be some lack of coherence in some places, or occasional unnatural expressions. The conversation can generally stay on topic but still has room for improvement to better simulate natural language communication.
4 points: fairly natural. The conversation is generally fluent and can convey meaning and emotions well. Although there may be occasional minor unnatural aspects, overall, it closely resembles real human dialogue. The conversation is coherent, able to closely follow the topic, and demonstrates good adaptability and understanding.
5 points: very natural. The conversation is extremely fluent and natural as if it were a real interaction with a person. There are no language or logical inconsistencies throughout the conversation, maintaining consistency and relevance to the topic. The expression is precise, and adaptable, closely simulating human communication habits and emotional responses, giving a very authentic and comfortable feeling.
Table 10: Prompt for evaluating the dialogue naturalness.